# On the Ability of Deep Networks to Learn Symmetries from Data: A Neural Kernel Theory

### v1

@@ 3.5 Extension to equivariant architectures
Through these examples, we build a framework that should be easily expandable to characterize how equivariant architectures and dataset symmetries interact in general.
@@ 3.5.1 Definitions
Dataset with translational symmetry.
$\mathcal{O}_{T}$ =$=\{g_{T}^{0}.x_{s},g_{T}^{1}.x_{s},\cdots g_{T}^{n-1}.x_{s}\},$ (13)
@@
$g_{T}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}(i_{x}+1)\bmod n\\ i_{y}\end{bmatrix}\right)$ (14)
Dataset with rotational symmetry.
$\mathcal{O}_{R}$ =$=\{g_{R}^{0}.x_{s},g_{R}^{1}.x_{s},g_{R}^{2}.x_{s},g_{R}^{3}.x_{s}\}$ (15)
@@
$g_{R}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}i_{y}\\ n-i_{x}\end{bmatrix}\right)$ (16)
Fully connected convolutional network (FC).
$f_{\text{FC}}(x)=A\frac{1}{\sqrt{k}}\phi(B\circledast x)_{v},$ (17)
@@
Global Average Pooling convolutional network (GAP).
We consider a convolutional network with global average pooling at the last layer. This network is invariant to discrete translations. The network $f_{\text{GAP}}:\mathbb{R}^{n\times n}\to\mathbb{R}$ is parameterized by:
$f_{\text{GAP}}(x)=\frac{1}{\sqrt{k}n^{2}}\sum_{k}\sum_{i_{x}}\sum_{i_{y}}A_{1k}\phi(B_{k,i_{x},i_{y}}\cdot x)$ (18)
where $A\in\mathbb{R}^{1\times k}$, and $B\in\mathbb{R}^{k\times 1\times 3\times 3}$ with $B_{k}\in\mathbb{R}^{1\times 1\times 3\times 3}$ indexing filter $k$ of $B$. $B_{k,i_{x},i_{y}}\in\mathbb{R}^{1\times 1\times n\times n}$ is obtained by centering $B_{k}$ at coordinates $(i_{x},i_{y})$ of an $n\times n$ grid with periodic boundary conditions, and filling the remaining entries with zeros. Then, the dot product is understood as the sum of the elementwise multiplications of all entries of $x$ and $B_{k,i_{x},i_{y}}$. We remark that this operation is effectively an alternative way of describing a convolution with filter bank $B$, but this indexing choice proves useful in the proofs (see App. C). After applying a convolutional layer to the data, this network averages the resulting representation across each of the $k$ output channels, and then takes a linear combination of these averages using a fully connected layer.
3.5.2 Theoretical results
@@
Convolutional neural networks, however, are not dot product kernels, as pixel proximity plays a role in the kernel similarity between two images (Arora et al.,, 2019). Nevertheless, here we show that the kernel matrix of convolutional neural networks is also circulant over common representations of the cyclic group, regardless of whether the equivariance matches the symmetry of the data or not. We thus get that the same spectral theory of generalization applies to these equivariant architectures. Furthermore, a well-known special case arises when the network is designed to be fully invariant to the symmetry of interest, which is also well captured by our spectral theory.
@@
See App. C for a proof of this proposition and the propositions below.
The kernel matrix of a fully connected convolutional network being circulant over a translation orbit, it can be analyzed in the Fourier domain, as done in the previous sections for a MLP. The same spectral formula for generalization error can thus be derived for this architecture and symmetry. Moreover, the kernel matrix is in general not rank-deficient the translation orbit. This implies that, a priori, no inverse eigenvalue will diverge in the denominator of Eq. 7. Were this to happen, the spectral error would go to 0, i.e., the network would achieve perfect generalization. The same conclusion about the inability of a MLP to extrapolate symmetries to partially observed classes thus holds for a fully connected convolutional network on a dataset with translational symmetry.
Proposition 2.
@@
The kernel matrix of a global average pooling convolutional network is not only circulant, but also constant. The kernel matrix computed over a dataset made of two orbits of two different classes is thus blockwise-constant in 2x2 blocks, one for each orbit, and by consequence rank-deficient. This ensures that some of the eigenvalues of the kernel matrix are null, leading the denominator of the spectral error (Eq. 7) to diverge and thus perfect generalization (0 error)555provided that the last frequency has non-zero power, i.e., the two classes are not fully collapsed in kernel space.. We thus recover the well-known fact that a convolutional network with global average pooling at the last layer makes predictions that are by construction invariant to translations, and thus correctly generalizes from a partial view of a class orbit.
@@
For a fully connected convolutional network applied to a dataset with rotational symmetry, the kernel matrix over an orbit remains circulant. Therefore, the same spectral theory of generalization applies as before. The kernel matrix is in general neither constant nor rank-deficient, and thus generalization on the missing point is not guaranteed to succeed.
Proposition 4.
@@
3.5.3 Empirical results
@@
Figure 7: Analysis of the prediction behavior of a fully connected convolutional network on pairs of orbits from rotated-MNIST. Same caption as in Fig. 4.
Fully connected (FC) convolutional network on rotated-MNIST.
We first analyze the behavior of a fully connected (FC) convolutional network, on the task of classifying a missing point from a dataset composed of two orbits, each generated from 8 rotations of a seed MNIST digit. The FC kernel matrix over a rotation orbit is circulant, as predicted by Prop. 3, allowing us to apply the spectral formula as we did for a MLP. The spectral error predicts well the symmetrized NTK error, across all pairs of orbits tested (Fig. 5C). As with a MLP, the numerator and denominator values of the spectral error are approximately proportional to the distance between classes in input space and their average orbit density respectively (Fig.  7), allowing the same interpretation of the factors that lead to successful generalization as for the MLP.
Global average pooling (GAP) convolutional network on rotated-MNIST.
We next consider a global average pooling convolutional network, on the same task and datasets. Its kernel matrix over a rotation orbit is also circulant, as predicted by Prop. 4, allowing us to apply the spectral theory. The prediction errors are in general remarkably low (but, crucially, not 0) for this architecture on these datasets (Fig. 5D). While we lack a comprehensive theory for this success, we speculate that the GAP convolutional network finds similarities in the local structure of images belonging to the same rotation orbit, allowing it to outperform other architectures. Interestingly, we find that a GAP convolutional network architecture where the filters don’t overlap (i.e., the stride matches filter size) fails to correctly classify all pairs of orbits (App. Fig. 16).
@@
Finally, as shown in App. G.1, the numerator and denominator values of the spectral error do not map well to the distance between classes in input space and their average orbit density. This illustrates the important fact that, for some neural architectures, the spectral quantities computed in kernel space are not direct correlates of simple geometric quantities in input space, and cannot be interpreted as simply.
Convolutional networks on translated-MNIST.
We next analyze the behavior of convolutional networks on pairs of orbits of translated-MNIST, a version of MNIST where the digits are translated along the image x-axis, with periodic boundary condition (App. G.2). Using a fully connected convolutional network, we do not see good generalization, as predicted by Prop. 1, and in good quantitative agreement with the spectral formula. Using a global average pooling convolutional network, we see perfect generalization of the network on the leave-out class (error is 0 to numerical precision), as expected from the fact that this architecture is invariant to translation. In our spectral formula, this perfect generalization is realized by the fact that the denominator of Eq. 7 diverges, as a consequence of the kernel matrix being rank-deficient (a consequence of Prop. 2).
@@ 4 Discussion
In this work we establish a theory on when, and to what extent, deep networks are capable of learning symmetries from data. We find that the generalization behavior of conventional networks trained with supervision on datasets presenting a cyclic-group symmetry is captured by a simple ratio of inverse kernel frequency powers. Our analysis of this formula sheds light on the limitations of conventional architectures trained with supervision to learn symmetries from data. In particular, we find that they are generally unable to extrapolate symmetries observed exhaustively on some classes to other partially sampled classes. Accurate generalization is only possible when the local structure of the data in kernel space (determined by network architecture) allows for correct generalization. In other words, conventional networks have no mechanism to learn symmetries that have not been embedded in their architecture a priori through equivariance.
@@
This result may seem at odds with the notion that deep networks are biased to find simple solutions (Zhang et al.,, 2021; Shah et al.,, 2020; Power et al.,, 2022; Humayun et al.,, 2024; Ortiz-Jiménez et al.,, 2020). In our study, networks are blind to the simple symmetric structure of the data. Were they to notice this simple structure, they would achieve data-efficient generalization. However, results showing that networks have a bias towards simplicity typically define simplicity as a notion of local structure in the data. These studies show that networks find the smoothest possible function capturing the training set. Here, however, the symmetric structure of the data is not a local structure: rotated versions of a digit are not close in pixel space to the upright digit, and they may even be closer to a different digit class altogether. For example, a ‘4’ upside down might look more like a ‘6’ than like another upright ‘4’. It is thus expected that networks with a bias for smooth solutions will fail on such task, in the absence of a mechanism to detect symmetries. Networks biased to find the shortest description of a dataset (in the sense of Kolmogorov complexity) may perform better on this task (Valle-Perez et al.,, 2019), but progress in this area is limited.
@@
On the other hand, our findings are compatible with the scaling laws observed when training deep networks (Kaplan et al.,, 2020; Bahri et al.,, 2024). Scaling laws show that deep networks continuously improve on generalization as the number of samples in the training set increases. Mapped to our setup, this corresponds to our observation that network performance gradually increases with the number of angles sampled from the group orbits. We see no abrupt change in generalization performance that would indicate that the network captures or “groks” the symmetry invariance, even with very long training times. An important implication of this observation is that networks should be data-inefficient on data presenting symmetries, as confirmed empirically by the ever-growing datasets used for training them (e.g., in computer vision JFT-300M (Sun et al.,, 2017), IG-3.6B (Singh et al.,, 2022) and LAION-5B (Schuhmann et al.,, 2022)). As the number of symmetric transformations present in the dataset increases (e.g., image rotation, translation, scaling, etc.), we only expect this generalization difficulty to grow, as this leads to a combinatorial explosion of possible transformations (Schott et al.,, 2021). This combinatorial explosion could partly be responsible for the brittleness of current deep learning approaches on edge cases, as observed for instance in Abbas and Deny, (2023).
@@
Many methods have been proposed to learn symmetries from data. For example, some propose to leverage more supervision by pairing images undergoing a symmetric transformation during training. This is the case, for instance, of self-supervised learning approaches via joint embedding (Geiping et al.,, 2023), and of autoencoder approaches (Dupont et al.,, 2020; Connor and Rozell,, 2020; Connor et al.,, 2024; Keller and Welling, 2021b, ), the goal of which is to map one sample to a transformed version of itself via an autoencoder equipped with transformation operators in latent space. Empirically, these methods are found to somewhat generalize group transformations (e.g., 3D rotations) to object classes that were not seen during training. Other directions include meta-learning approaches (Finn et al.,, 2017; Yang and Hu,, 2020), feature learning approaches (Yang et al.,, 2023) and dynamic architecture choices (Chauhan et al.,, 2024; Stanley and Miikkulainen,, 2002), which could potentially alleviate the problem of having a fixed, frozen, kernel incapable of adapting to problem symmetries. Indeed, our understanding is that networks cannot learn symmetries because their kernel is defined by their architecture, and cannot adapt to the symmetries of a given dataset. Designing mechanisms to adapt the kernel induced by the network to the symmetries of the problem at stake seems a likely path forward to devise more data-efficient deep networks. In future work, we would like to extend our theory to these methods, and establish whether they are theoretically able to learn symmetries from data, and if so, to what extent.
@@
Our theory could also be extended in various ways to capture a richer phenomenology. First, we have focused on the simple discrete one-dimensional cyclic group in this study. It would be interesting to understand whether and how the theory could be extended to continuous, high-dimensional and more complicated groups, such as SO(3). Second, our study focuses on symmetries that are “native” to the space in which the datasets reside (e.g., image rotations). However, group actions may more realistically exist in a latent space affecting the dataset indirectly (e.g., images of 3D-rotated objects). This latter case is particularly interesting because architectures cannot easily be designed to be equivariant to symmetries which are not directly acting in dataset space. We note, however, that given the failure of conventional architectures to correctly generalize simple symmetries native to the dataset space, we do not expect these same architectures to be able to learn more complex, latent space symmetries.
@@
A source of inspiration for learning symmetries could be found in cognitive science and neuroscience. Empirically, there is evidence that humans are superior to deep networks at some tasks necessitating symmetry invariance. Recently, Ollikka et al., (2024) show that humans beat state-of-the-art deep networks and most vision-language models at recognizing objects in unusual poses. Humans are also known to be able to reason about problem symmetries, for example in the problem of mental rotation (Shepard and Metzler,, 1971). Interestingly, the time for subjects to compare one 3D object to another in mental rotation tasks is proportional to the angular difference between the two objects, implying that some recurrent processes in the brain may be involved, recurrence lacking from current state-of-the art architectures in deep learning. There is also evidence that the mental rotation ability is learned through experience (or at least not fully operational at birth), as studies on infants show that there is a critical age when they are able to perform this task (Bambha et al.,, 2022). Conversely, children learning to read need to first unlearn mirror symmetries in order to differentiate characters from their mirrored version (Perea et al.,, 2011; Pegado et al.,, 2011). Abilities to capture specific problem symmetries may, however, be partially or fully innate in some animals. A study on chicks show that they have the ability to recognize an unknown synthetic 3D object from birth, across multiple points of view (Wood,, 2013). This finding prompts us to reconsider whether humans and animals actively learn problem symmetries, or rely on neural architectures which are pre-configured for specific symmetries. It is at least clear from behavioral experiments that humans have the ability to adapt to unnatural symmetric transformations of their environment. Classically, Kohler and Erismann (Kohler,, 1963) showed that a subject wearing goggles reversing the world upside down can adapt to this new environment after a few days of practice. The extent to which this adaptation is possible for new, previously unseen symmetries is unclear, however (e.g., what if the goggles permanently permuted the location of every pixel?). Finally, a direction of interest may be to imitate the physicist: by describing the world symbolically, the physicist comes to discover the symmetries of the world and exploit them for predictions. Language models and other neuro-symbolic approaches may be able to partly replicate this ability, or at least piggy-back on human accumulated knowledge about the symmetries of the world.
Acknowledgments
We thank David Radnell, Milo Orlich, Samuel A. Ocko and Luigi Acerbi for useful discussions, Martin Trapp for comments on the manuscript, and Adityanarayanan Radhakrishnan for sharing with us his clear and concise lecture notes on neural kernels.

### v2

@@ 4.6 Extension to Equivariant Architectures
Through these examples, we build a framework that should be easily extensible to characterize how equivariant architectures and dataset symmetries interact in general.
@@ 4.6.1 Definitions
Definition 7 (Dataset with translational symmetry)
A dataset with translational symmetry is composed of seed images and all their translations. For the purpose of the proofs to follow, we will focus on a single orbit of this dataset, which consists of a single seed point $x_{s}\in\mathbb{R}^{n\times n}$ and all of its translations:
$\mathcal{O}_{T}$ =$=\{g_{T}^{0}.x_{s},g_{T}^{1}.x_{s},\cdots g_{T}^{n-1}.x_{s}\},$
@@
$g_{T}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}(i_{x}+1)\bmod n\\ i_{y}\end{bmatrix}\right)$
Definition 8 (Dataset with rotational symmetry)
A dataset with rotational symmetry is composed of seed images and all their rotations in $C_{4}$ (we limit ourselves to 4 cardinal rotations to avoid definitional problems of image rotation on discrete pixel grids). We will focus on a single orbit of this dataset, which consists in a single seed point $x_{s}\in\mathbb{R}^{n\times n}$ and all its rotations:
$\mathcal{O}_{R}$ =$=\{g_{R}^{0}.x_{s},g_{R}^{1}.x_{s},g_{R}^{2}.x_{s},g_{R}^{3}.x_{s}\}$
@@
$g_{R}.x\left(\begin{bmatrix}i_{x}\\ i_{y}\end{bmatrix}\right)=x\left(\begin{bmatrix}i_{y}\\ n-i_{x}\end{bmatrix}\right)$
Definition 9 (Fully connected convolutional network (FC))
A fully connected convolutional network is a network $f:\mathbb{R}^{n\times n}\to\mathbb{R}$ parameterized by:
$f_{\text{FC}}(x)=A\frac{1}{\sqrt{k}}\phi(B\circledast x)_{v},$
@@
Definition 10 (Global Average Pooling convolutional network (GAP))
A GAP network is a convolutional network with global average pooling at the last layer. This network is invariant to discrete translations. The network $f_{\text{GAP}}:\mathbb{R}^{n\times n}\to\mathbb{R}$ is parameterized by:
$f_{\text{GAP}}(x)=\frac{1}{\sqrt{k}n^{2}}\sum_{k}\sum_{i_{x}}\sum_{i_{y}}A_{1k}\phi(B_{k,i_{x},i_{y}}\cdot x)$
4.6.2 Theoretical Results on Equivariant Architectures
@@
Convolutional neural networks, however, are not dot product kernels, as pixel proximity plays a role in the kernel similarity between two images (Arora et al., 2019). Nevertheless, here we show that the kernel matrix of convolutional neural networks is also circulant over common representations of the cyclic group, regardless of whether the equivariance matches the symmetry of the data or not. We thus get that the same spectral theory of generalization applies to these equivariant architectures. Furthermore, a well-known special case arises when the network is designed to be fully invariant to the symmetry of interest, which is also well captured by our spectral theory.
@@
See App. E for a proof of this proposition and the propositions below.
The kernel matrix of a fully connected convolutional network being circulant over a translation orbit, it can be analyzed in the Fourier domain, as done in the previous sections for a MLP. The same spectral formula for generalization error can thus be derived for this architecture and symmetry. Moreover, the kernel matrix is in general not rank-deficient on the translation orbit. This implies that, a priori, no inverse eigenvalue will diverge in the denominator of Eq. 1. Were this to happen, the spectral error would go to 0, i.e., the network would achieve perfect generalization. The same conclusion about the inability of a MLP to extrapolate symmetries to partially observed classes thus holds for a fully connected convolutional network on a dataset with translational symmetry.
Proposition 12
@@
The kernel matrix of a global average pooling convolutional network is not only circulant, but also constant. The kernel matrix computed over a dataset made of two orbits of two different classes is thus blockwise-constant in 2x2 blocks, one for each orbit, and by consequence rank-deficient. This rank-deficiency ensures that some of the eigenvalues of the kernel matrix are null, leading the denominator of the spectral error (Eq. 1) to diverge and thus perfect generalization (0 error). [^8] This is true in the general case where the last frequency (at the numerator of the spectral formula) has non-zero power, i.e., the two classes are not fully collapsed in kernel space. We thus recover the well-known fact that a convolutional network with global average pooling at the last layer makes predictions that are by construction invariant to translations, and thus correctly generalizes from a partial view of a class orbit.
@@
For a fully connected convolutional network applied to a dataset with rotational symmetry, the kernel matrix over an orbit remains circulant. Therefore, the same spectral theory of generalization applies as before. The kernel matrix is in general neither constant nor rank-deficient, and thus generalization on the missing point is not guaranteed to succeed. [^9] This result, however, does not preclude a neural classifier from classifying the point correctly. Indeed, if the error is less than 1, a classifier would classify the point correctly (since the prediction is closer to the correct label than the other label which is -1 in our setup). Whether the neural classifier will classify the point correctly depends on the specific interplay between orbit density and class separation in kernel space, which depends on geometric properties of the data and of the architecture, as described by the spectral formula.
Proposition 14
@@
4.6.3 Empirical Results on Equivariant Architectures
@@
Figure 7: Analysis of the prediction behavior of a fully connected convolutional network on pairs of orbits from rotated-MNIST. Same caption as Fig. 4.
Fully connected (FC) convolutional network on rotated-MNIST
We first analyze the behavior of a fully connected (FC) convolutional network, on the task of classifying a missing point from a dataset composed of two orbits, each generated from 8 rotations of a seed MNIST digit. The FC kernel matrix over a rotation orbit is circulant, as predicted by Prop. 3, allowing us to apply the spectral formula as we did for a MLP. The spectral error predicts well the symmetrized NTK error, across all pairs of orbits tested (Fig. 5D). As with a MLP, the numerator and denominator values of the spectral error are approximately proportional to the distance between classes in input space and their average orbit density respectively (Fig.  7), allowing the same interpretation of the factors that lead to successful generalization as for the MLP.
Global average pooling (GAP) convolutional network on rotated-MNIST
We next consider a global average pooling convolutional network, on the same task and datasets. Its kernel matrix over a rotation orbit is also circulant, as predicted by Prop. 4, allowing us to apply the spectral theory. The prediction errors are in general remarkably low (but, crucially, not 0) for this architecture on these datasets (Fig. 5E).
While we lack a comprehensive theory for this success, we speculate that the GAP convolutional network finds similarities in the local structure of images belonging to the same rotation orbit, allowing it to outperform other architectures. Interestingly, we find that a GAP convolutional network architecture where the filters don’t overlap (i.e., the stride matches filter size) fails to correctly classify some pairs of orbits (App. Fig. 16).
@@
Finally, as shown in App. I.1, the numerator and denominator values of the spectral error do not map well to the distance between classes in input space and their average orbit density. This illustrates the important fact that, for some neural architectures, the spectral quantities computed in kernel space are not direct correlates of simple geometric quantities in input space, and cannot be interpreted as simply. Nevertheless, even in these cases, the error on a missing point is given by the spectral formula (with the approximations mentioned), where the quantities in the formula represent geometric quantities in kernel space.
Convolutional networks on translated-MNIST
We next analyze the behavior of convolutional networks on pairs of orbits of translated-MNIST, a version of MNIST where the digits are translated along the image x-axis, with periodic boundary condition (App. I.2). Using a fully connected convolutional network, we do not see good generalization, as predicted by Prop. 1, and in good quantitative agreement with the spectral formula. Using a global average pooling convolutional network, we see perfect generalization of the network on the leave-out class (error is 0 to numerical precision), as expected from the fact that this architecture is invariant to translation. In our spectral formula, this perfect generalization is realized by the fact that the denominator of Eq. 1 diverges, as a consequence of the kernel matrix being rank-deficient (a consequence of Prop. 2).
In conclusion, we find that the generalization behavior of convolutional architectures on symmetric datasets is governed by the same formula and depends on the same geometric quantities as in the MLP case (namely orbit separation and density in kernel space). For highly non-linear architectures such as networks with a global-average pooling at the last layer, it is important to note that the geometric quantities in kernel space entering our spectral error are no longer trivially correlated to the same geometric quantities in input space, making the interpretation of these quantities kernel-specific. Finally, we find that only when the network is designed to be fully invariant to the symmetry of interest (e.g., GAP-CNN for translations), can it be theoretically guaranteed that the network will systematically generalize the symmetry correctly to unseen classes.
@@ 5 Discussion
In this work we establish a theory on when—and to what extent—deep networks are capable of learning symmetries from data. We find that the generalization behavior of conventional networks trained with supervision on datasets presenting a cyclic-group symmetry is captured by a simple ratio of inverse kernel frequency powers. Our analysis of this formula sheds light on the limitations of conventional architectures trained with supervision to learn symmetries from data. In particular, we find that they are generally unable to extrapolate symmetries observed exhaustively on some classes to other partially sampled classes. Accurate generalization is only possible when the local structure of the data in kernel space (determined by network architecture) allows for correct generalization. In other words, conventional networks have no mechanism to learn symmetries that have not been embedded in their architecture a priori through equivariance.
What are the practical implications of our results? There may be different remedies to the issues described above, which may also depend on the specific constraints of the problem. We outline here these solutions briefly before entering a more thorough discussion. First, resorting to very large datasets or having prior knowledge about the symmetries of the problem so as to build them into the model via data augmentation or equivariance may suffice in some cases. When this is not possible, introducing soft biases encouraging the network to find equivariant structures may be a fruitful alternative. Pretraining with self-supervised learning losses using the symmetries of interest as augmentation may also alleviate the problem of generalization, but it would require further studies to understand how such pretraining affects learned representations. We cannot exclude that in the feature-rich learning regime, symmetries can be learned and generalized—at least in some cases. Finally, there may be some entirely new learning procedures which could somehow respect the symmetries of the data without explicitly building these symmetries into the model.
Could neural networks learn symmetries outside the NTK regime? Our theory is only valid in the NTK regime, also sometimes referred to as the ‘lazy regime’, which assumes (1) infinite width, and (2) a certain scaling of the parameters with width (weights scale as $1/\sqrt{N}$ where $N$ is layer width). We found in our experiments that the theory also recapitulates well the training dynamics of finite-width networks using Pytorch default scaling of weights (Kaiming Uniform, weights scale as $1/\sqrt{N}$). However, we cannot exclude that for other initialization schemes, datasets (e.g., CIFAR-10, ImageNet) or network architectures (e.g., deeper MLPs), finite-width networks could successfully learn the symmetry of the problem. Furthermore, other regimes of weight scaling have been studied in the infinite limit, for example the ‘mean-field’ regime (where weights scale as $1/N$) (Bach, 2017; Mei et al., 2018; Chizat and Bach, 2018), the ‘high-dimensional’ regime (Saglietti et al., 2022), and the ‘dynamical mean-field’ regime (Bordelon and Pehlevan, 2022; Yang et al., 2024a). These regimes, often referred to as ‘feature learning’ or ‘active’ regimes, are characterized by a kernel that evolves in time, and thus learns features. This leads to different generalization behaviors both in the finite-width (Geiger et al., 2020) and infinite-width limit (Bordelon and Pehlevan, 2022). It would be interesting to study symmetry learning in these regimes. In a recent study, Jacot et al. (2025) showed that deep networks in the active regime are able to learn data symmetries. However, their definition of symmetry is that of an invariant subspace warped by a non-linear function (through low-index functions, i.e., functions of the form $f(x)=g(Ax)$ where $A$ is a low rank matrix). In contrast, we define a symmetry as the non-local, linear action of a group on a dataset.
@@
Our results may seem at odds with the notion that deep networks are biased to find simple solutions (Shah et al., 2020; Ortiz-Jimenez et al., 2020; Zhang et al., 2021; Power et al., 2022; Humayun et al., 2024). In our study, networks are blind to the simple symmetric structure of the data. Were they to notice this simple structure, they could achieve data-efficient generalization. Importantly, results showing that networks have a bias towards simplicity typically define simplicity as a notion of local structure in the data. These studies show that networks find the smoothest possible function capturing the training set. Here, however, the symmetric structure of the data is not a local structure: rotated versions of a digit are not close in pixel space to the upright digit, and they may even be closer to a different digit class altogether. For example, a ‘4’ upside down might look more like a ‘6’ than like another upright ‘4’. It is thus expected that networks with a bias for smooth solutions will fail on such task, in the absence of a mechanism to detect symmetries. Networks biased to find the shortest description of a dataset (in the sense of Kolmogorov complexity) may perform better on this task (Valle-Perez et al., 2019), but progress in this area is limited. Bayesian model selection approaches may be a promising avenue to find such shortest description (van der Wilk et al., 2018; Immer et al., 2022; van der Ouderaa and van der Wilk, 2022; van der Ouderaa et al., 2023), and have recently been applied to learning data symmetries (van der Ouderaa et al., 2024).
@@
On the other hand, our findings are compatible with the scaling laws observed when training deep networks (Kaplan et al., 2020; Bahri et al., 2024). Scaling laws show that deep networks continuously improve on generalization as the number of samples in the training set increases. Mapped to our setup, this corresponds to our observation that network performance gradually increases with the number of angles sampled from the group orbits. We see no abrupt change in generalization performance that would indicate that the network captures or “groks” the symmetry invariance, even with very long training times. The absence of a mechanism to capture data symmetries could play a role in the notable data-inefficiency of current deep learning approaches, inefficiency demonstrated by the evergrowing datasets used for training them (e.g., in computer vision JFT-300M (Sun et al., 2017), IG-3.6B (Singh et al., 2022) and LAION-5B (Schuhmann et al., 2022)). As the number of symmetric transformations present in the data increases (e.g., image rotation, translation, scaling, etc.), we only expect the generalization difficulty to increase, as this leads to a combinatorial explosion of possible transformations (Schott et al., 2022). This combinatorial explosion could partly be responsible for the brittleness of current deep learning approaches on edge cases, as observed for instance in Abbas and Deny (2023), where it is shown that deep networks for vision fare particularly poorly on combinations of symmetric transformations such as scaling and rotation.
@@
Many methods have been proposed to learn symmetries from data. For example, some propose to leverage more supervision by pairing images undergoing a symmetric transformation during training. This is the case, for instance, of self-supervised learning approaches via joint embedding (Balestriero et al., 2023), and of autoencoder approaches (Dupont et al., 2020; Connor and Rozell, 2020; Keller and Welling, 2021a; Connor et al., 2024), the goal of which is to map one sample to a transformed version of itself via an autoencoder equipped with transformation operators in latent space. Empirically, these methods are found to somewhat generalize group transformations (e.g., 3D rotations) to object classes that were not seen during training. Other directions include meta-learning approaches (Finn et al., 2017; Yang and Hu, 2021), feature learning approaches (Bach, 2017; Yang et al., 2024a; Jacot et al., 2025) and dynamic architecture choices (Stanley and Miikkulainen, 2002; Chauhan et al., 2024), which could potentially alleviate the problem of having a fixed, frozen, kernel incapable of adapting to problem symmetries. Indeed, our understanding is that networks cannot learn symmetries because their kernel is defined by their architecture, and cannot adapt to the symmetries of a given dataset. Designing mechanisms to adapt the kernel induced by the network to the symmetries of the problem at stake seems a likely path forward to devise more data-efficient deep networks. In future work, our theory could be extended to these methods, in order to establish whether they are theoretically able to learn symmetries from data, and if so, to what extent.
@@
Our theory could also be extended in various ways to capture a richer phenomenology. First, we have mainly focused here on the simple discrete one-dimensional cyclic group. However, our theoretical framework is readily extendable to any finite group including non-abelian groups (see Theorem 15), and it would be interesting to study the practical implications of the theory for such more complex groups. It would also be interesting to understand whether and how the theory could be extended to continuous groups, such as $SO(3)$. Second, our study focuses on symmetries that are native to the space in which the datasets reside (e.g., image rotations). However, group actions may more realistically exist in a latent space affecting the dataset indirectly (e.g., images of 3D-rotated objects). This latter case is particularly interesting because architectures cannot easily be designed to be equivariant to symmetries which are not directly acting in dataset space. We note, however, that given the failure of conventional architectures to correctly generalize simple symmetries native to the dataset space, we do not expect these same architectures to be able to learn more complex, latent space symmetries.
@@
A source of inspiration for learning symmetries could be found in cognitive science and neuroscience. Empirically, there is evidence that humans are superior to deep networks at some tasks necessitating symmetry invariance. In a recent example, Ollikka et al. (2025) showed that humans beat state-of-the-art deep networks and most vision-language models at recognizing objects in unusual poses. Humans are also known to be able to reason about problem symmetries, for example in the problem of mental rotation (Shepard and Metzler, 1971). Interestingly, the time for subjects to compare one 3D object to another in mental rotation tasks is proportional to the angular difference between the two objects, implying that some recurrent processes in the brain may be involved, recurrence lacking from current state-of-the art architectures in deep learning. There is also evidence that the mental rotation ability is learned through experience (or at least not fully operational at birth), as studies on infants show that there is a critical age when they are able to perform this task (Bambha et al., 2022). Conversely, children learning to read need to first unlearn mirror symmetries in order to differentiate characters from their mirrored version (Perea et al., 2011; Pegado et al., 2011). The ability to capture specific problem symmetries may, however, be partially or fully innate in some animals. A study on chicks show that they have the ability to recognize an unknown synthetic 3D object from birth, across multiple points of view (Wood, 2013). This finding prompts us to reconsider whether humans and animals actively learn problem symmetries, or rely on neural architectures which are pre-configured for specific symmetries. It is at least clear from behavioral experiments that humans have the ability to adapt to unnatural symmetric transformations of their environment. Classically, Kohler and Erismann (Kohler, 1963) showed that a subject wearing goggles reversing the world upside down can adapt to this new environment after a few days of practice. The extent to which this adaptation is possible for new, previously unseen symmetries is unclear, however (e.g., what if the goggles permanently permuted the location of every pixel?). Finally, a direction of interest may be to imitate the physicist: by describing the world symbolically, the physicist comes to discover the symmetries of the world and exploit them for predictions. Language models and other neuro-symbolic approaches may be able to partly replicate this ability, or at least piggy-back on human accumulated knowledge about the symmetries of the world.
Acknowledgments and Disclosure of Funding
We thank David Radnell, Milo Orlich, Samuel A. Ocko, Jaakko Lehtinen, Tycho F. A. van der Ouderaa, Arthur Jacot, Jan Gerken, Stefano Sarao Mannelli for useful discussions and additional references, Luigi Acerbi and Martin Trapp for comments on the manuscript, Adityanarayanan Radhakrishnan for his clear and concise lecture notes on neural kernels, and Fabio Anselmi for an insight that led to the proof of Theorem 15. We thank the anonymous reviewers and our editor Aapo Hyvärinen for their constructive feedback and suggestions, and for their time. Funding: Research Council of Finland grant to S.D. under the Project “Neuroscience-inspired Deep Learning”: 3357590.